The NetBSD Project

Chris Demetriou January 18, 1995

What is NetBSD?

• A free and complete BSD operating system

• Derived from 4.4BSD-Lite and Berkeley’s Net/2 release • Runs on many hardware platforms

• Distributed with complete source

• Supported by volunteers

Some of the NetBSD Project’s Major Goals:

• Provide a free, complete, stable BSD system

• “Support” for popular platforms

• Avoid encumbering licenses Summary:

Provide a free system that professionals and hobbyists can use for whatever they want.

“User-land Improvements”

• Continuing 4.4BSD-Lite merge

• Weak references, for library symbols

• Updated some of the third-party software

• Getting closer to POSIX.2 and other standards • Many, many bug fixes and “little” improvements

Kernel Improvements

• More architectures supported

• 4.4BSD-Lite file system code

• Union file system

• Most of the system is now 64-bit clean • Ever closer to POSIX.1 conformance

More Kernel Improvements

• SCSI subsystem seriously cleaned up and improved • Most ports now use new autoconfig mechanism

• More work done on the diskless boot code

• Performance improvements for several ports

• And many, many more “little” fixes and improvements

The NetBSD Project, January 18, 1995 6

Working ports:

i386/i486/Pentium  Amiga
HP9000/300  DECstation
Macintosh (some models) Sun3
SPARC (sun4, sun4c)  pc532 (ns32k board)

Ports in progress:

Alpha ARM IBM RT PC PowerPC VAX

Ports we want to see done:

SPARC (sun4m) HP9000/700

Want a port? Support one!

What you need:

• One machine to start, usually a second later • Technical reference material

• A lot of dedication, patience, and time We’ll help, to the extent that we can. Best way to get a port done:

“JUST DO IT!”

Compatibility with Other Systems

• BSDI binary compatibility (i386)

• SunOS binary compatibility (SPARC, m68k)

• SVR4 binary compatibility (i386)

• Nearly 4.4BSD source-code compatible

• POSIX It’s easy to port software to NetBSD.

4.4BSD-Lite Integration

Goals:

• Integrate 4.4BSD-Lite changes.

• Roll forward NetBSD improvements. Status:

• Kernel done long ago.

• User-land mostly done. Many, many fixes and improvements sent back to Berkeley, for integration into the final release of 4.4BSD-Lite.

Future work on NetBSD

• Finish 4.4-Lite integration

• Support more platforms

• Continue POSIX compliance work • Performance enhancements

• Improve RFC support

• Improve usability where possible • Fulfill user demands as best we can • Anything else we can think of…

NetBSD Releases

• NetBSD 0.8 – April, 1993

• NetBSD 0.9 – August, 1993

• NetBSD 1.0 – October, 1994

We will ship no software before its time.

Also available: • NetBSD-current – Daily snapshot of work in progress

NetBSD-current

A snapshot of our work in progress. • New source snapshot nightly.

• Source tar files created weekly

• Binary snapshots occasionally

• Can be obtained via anonymous ftp and SUP NetBSD-current helps people get our sources quicker, and

allows others to find and fix our bugs more easily!

Who does the work?

A “cast of thousands” of contributors including: • CSRG, for Net/2, 4.4-Lite, and for “BSD” in general • The NetBSD core team, port maintainers, and other

developers

• Users who have sent in bug reports, fixes, and enhancements

• The “NetBSD-current” alpha-testers How to get NetBSD via anonymous FTP:

    ftp.NetBSD.ORG:pub/NetBSD
    ftp.iastate.edu:pub/netbsd

NetBSD mailing lists:

    majordomo@NetBSD.ORG

And if you have questions or comments:

  • Send mail to netbsd-comments@NetBSD.ORG

• Come talk to us privately

• Ask questions now!